# Flash Diffusion: Accelerating Any Conditional Diffusion Model for Few Steps Image Generation

### v2

@@
Abstract
In this paper, we propose an efficient, fast, and versatile distillation method to accelerate the generation of pre-trained diffusion models: Flash Diffusion. The method reaches state-of-the-art performances in terms of FID and CLIP-Score for few steps image generation on the COCO2014 and COCO2017 datasets, while requiring only several GPU hours of training and fewer trainable parameters than existing methods. In addition to its efficiency, the versatility of the method is also exposed across several tasks such as text-to-image, inpainting, face-swapping, super-resolution and using different backbones such as UNet-based denoisers (SD1.5, SDXL) or DiT (Pixart-$\alpha$), as well as adapters. In all cases, the method allowed to reduce drastically the number of sampling steps while maintaining very high-quality image generation. The official implementation is available at https://github.com/gojasper/flash-diffusion.
# Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@ 1 Introduction
Diffusion Models (DM) [64, 17, 67] have proven to be one of the most efficient class of generative models for many tasks ranging from image synthesis [8, 55, 57, 47], video generation [19, 10, 3, 2], audio [29] or 3D [41, 40, 74]. They have raised particular interest and enthusiasm for text-to-image applications [54, 55, 57, 59, 18, 11, 50, 5, 6] where they outperform other approaches. However, their usability for real-time applications remains limited by the intrinsic iterative nature of their sampling mechanism. At inference time, these models aim at iteratively denoising a sample drawn from a Gaussian distribution to finally create a sample belonging to the data distribution. Nonetheless, such a denoising process requires multiple evaluations of a potentially very computationally costly neural function such that their faster counterparts such as Generative Adversarial Networks (GANs) [12] and Variational AutoEncoders (VAEs) [27] still tend to be preferred.
@@
Recently, more efficient solvers [37, 38, 80, 82] or diffusion distillation methods [60, 68, 33, 76, 36, 56, 43, 42, 61, 62, 77] aiming at reducing the number of sampling steps required to generate satisfying samples from a trained diffusion model have emerged to try to tackle this issue. Nonetheless, solvers typically require at least 10 Neural Function Evaluations (NFEs) to produce satisfying samples while distillation methods may require extensive training resources [36, 77, 45] or require an iterative training procedure to update the teacher model throughout training [60, 33, 30] limiting their applications and reach. Moreover, most of the existing distillation methods are tailored for a specific task such as text-to-image. It is still unclear how they would perform on other tasks, using different conditionings and diffusion model architectures. In addition, the most efficient approaches rely on an adversarial training procedure [76, 61, 33] potentially leading to unstable trainings and requiring extensive hyper-parameters tuning [62].
@@
In this paper, we present Flash Diffusion, a fast, robust, and versatile diffusion distillation method that allows to drastically reduce the number of sampling steps while maintaining a very high image generation quality. The proposed method aims at training a student model to predict in a single step a denoised multiple-step teacher prediction of a corrupted input sample. The method also drives the student distribution towards the real input sample manifold with an adversarial objective [12] and ensures that it does not drift too much from the learned teacher distribution using distribution matching [9, 31].
The method is compatible with LoRA [20] and is able to generate high-quality samples in only a few steps while requiring only several GPU hours of training and fewer trainable parameters than existing methods. We show that the method is able to reach State-of-the-Art (SOTA) performances in terms of FID and CLIP score for few steps image generation on the COCO2014 and COCO2017 datasets. In addition to its efficiency, the versatility of the method is also exposed across several tasks such as text-to-image, inpainting, face-swapping, image upscaling and using different diffusion models backbones (SD1.5 [57], SDXL [50] and Pixart-$\alpha$ [5]) as well as adapters [46]. In all cases, the method allows to reduce drastically the number of sampling steps while maintaining very high-quality image generation.
The main contributions of the paper are as follows:
@@
  • We validate the method for a text-to-image task and show that it is able to produce SOTA results for few steps image generation on standard benchmark datasets with only two NFEs, which is equivalent to a single step with classifier-free guidance while having far fewer training parameters than competitors and requiring only a few GPU hours of training.
@@
  • We emphasize the versatility of the method through an extensive experimental study across various tasks (text-to-image, image inpainting, super-resolution, face-swapping), diffusion model architectures (SD1.5, SDXL and Pixart-$\alpha$) and illustrate its compatibility with adapters [46].
  • Finally, we share our implementation that showcases the method’s reproducibility and stability, and potentially opens the door to a wider range of practical applications of Flash Diffusion. See https://github.com/gojasper/flash-diffusion.
2 Related Works
@@
Diffusion models consist in artificially corrupting input data according to a given noise schedule [64, 17, 67] such that the data distribution eventually resembles a standard Gaussian one. They are then trained to estimate the amount of noise added in order to learn a reverse diffusion process allowing them, once trained, to generate new samples from Gaussian noise. Those models can be conditioned with respect to various inputs such as images [57], depth maps, edges, poses [79, 46] or text [8, 55, 57, 47, 11, 18, 50] where they demonstrated very impressive results. However, the need to recourse to a large number of sampling steps (typically 50 steps) at inference time to generate high-quality samples has limited their application for real-time applications and narrowed their usability and reach.
@@
In order to tackle this limitation, several methods have recently emerged to reduce the number of function evaluations required at inference time. On the one hand, several papers tried to build more efficient solvers to speed up the generation process [37, 38, 80, 82] but these methods still require the use of several steps (typically 10) to generate satisfying samples. On the other hand, several approaches relying on model distillation [15] proposed to train a student network that would learn to match the samples generated by a teacher model but in fewer steps. A simple approach would consist in building pairs of noise/teacher samples and training a student model to match the teacher predictions from the same noise in a single step with a regression loss [39, 83]. Nonetheless, this approach remains quite limited and struggles to match the quality of the teacher model since there is no underlying useful information to be learned by the student in full noise. Building upon this idea, several methods were proposed to first apply the forward diffusion process to an input sample and then pass it to the student network. The student prediction is then compared to the learned distribution of the teacher model using either a regression loss [28, 77] an adversarial objective [76, 61, 62, 78] or distribution matching [77, 78].
@@
Progressive distillation [60, 45] is also a method that has proven to be quite promising. It consists in training a student model to predict a two-step teacher denoising of a noisy sample in a single step theoretically halving the number of required sampling steps. The teacher is then replaced by the new student and the process is repeated several times. This approach was also enriched with a GAN-based objective that allows to further reduce the number of sampling steps needed from 4-8 to a single pass [33]. InstaFlow [36] proposed instead to rely on rectified flows [35] to ease the one-step distillation process. However, this approach may require a significant number of training parameters and a long training procedure, making it computationally intensive.
@@
Consistency models [68, 65, 42, 24] is also a promising, effective, and one of the most versatile distillation methods proposed in the literature. The main idea is to train a model to map any point lying on the Probability Flow Ordinary Differential Equation (PF-ODE) to its origin, theoretically unlocking single-step generation. Luo et al. [43] combined Latent Consistency Model (LCM) and LoRAs [20] and showed that it is possible to train a strong student with a very limited number of trainable parameters and a few GPU hours of training. Nonetheless, those models still struggle to achieve single-step generation and reach the sampling quality of peers.
@@
In a parallel study conducted recently, the authors of [78] also introduced the combined use of a distribution matching loss and an adversarial loss, a method we also employ in our paper. Nonetheless, they do not rely on the use of a distillation loss that proved highly efficient in our experiments and do not compute the adversarial loss with respect to the same inputs. Moreover, their approach still necessitates training an other denoiser to assess the score of the fake samples, significantly increasing the number of trainable parameters and, consequently, the computational burden of the method. Furthermore, the ability of their method to generalize and perform effectively across different tasks and diffusion model architectures remains unclear.
3 Background
3.1 Diffusion Models
Let $x_{0}\in\mathcal{X}$ be a set of input data such that $x_{0}\sim p(x_{0})$ where $p(x_{0})$ is an unknown distribution. Diffusion models are a class of generative models that define a Markovian process $(x_{t})_{t\in[0,T]}$ consisting in creating a noisy version $x_{t}$ of $x_{0}$ by iteratively injecting Gaussian noise to the data $x_{0}$. This process is such that as $t$ increases the distribution of the noisy samples $x_{t}$ eventually becomes equivalent to an isotropic Gaussian distribution. The noise schedule is controlled by two differentiable functions $\alpha(t)$, $\sigma(t)$ for any $t\in[0,T]$ such that the log signal-to-noise ratio $\log[\alpha(t)^{2}/\sigma(t)^{2}]$ is decreasing over time. Given any $t\in[0,T]$, the distribution of the noisy samples given the input $q(x_{t}|x_{0})$ is called the forward process and is defined by $q(x_{t}|x_{0})=\mathcal{N}\left(x_{t};\alpha(t)\cdot x_{0},\sigma(t)^{2}\cdot\mathbf{I}\right)$ from which we can sample as follows:
$x_{t}=\alpha(t)\cdot x_{0}+\sigma(t)\cdot\varepsilon\hskip 10.00002pt\text{with}\hskip 10.00002pt\varepsilon\sim\mathcal{N}\left(0,\mathbf{I}\right)\,.$ (1)
The main idea of diffusion models is to learn to denoise a noisy sample $x_{t}\sim q(x_{t}|x_{0})$ in order to learn the reverse process allowing to ultimately create samples $\tilde{x}_{0}$ directly from pure noise. In practice, during training a diffusion model consists in learning a parametrized function $x_{\theta}$ conditioned on the timestep $t$ and taking as input the noisy sample $x_{t}$ such that it predicts a denoised version of the original sample $x_{0}$. The parameters $\theta$ are then learned via denoising score matching [69, 66].
@@ 3.1 Diffusion Models
$\mathcal{L}=\mathbb{E}_{x_{0}\sim p(x_{0}),t\sim\pi(t),\varepsilon\sim\mathcal{N}\left(0,\mathbf{I}\right)}\left[\lambda(t)\left\|x_{\theta}(x_{t},t)-x_{0}\right\|^{2}\right]\,,$ (2)
where $\lambda(t)$ is a scaling factor that depends on the timestep $t\in[0,1]$ and $\pi(t)$ is a distribution over the timesteps. Note that Eq. (2) is actually equivalent to learning a function $\varepsilon_{\theta}$ estimating the amount of noise $\varepsilon$ added to the original sample using the repametrization $\varepsilon_{\theta}(x_{t},t)=\big{(}x_{t}-\alpha(t)\cdot x_{\theta}(x_{t},t)\big{)}/\sigma(t)$. Song et al. [67] showed that $\varepsilon_{\theta}$ can be used to generate new data points from Gaussian noise by solving the following PF-ODE [67, 60, 25, 37]:
$\mathrm{d}x_{t}=\left[f(x_{t},t)-\frac{1}{2}g^{2}(t)\nabla\log p_{\theta}(x_{t})\right]\mathrm{d}t\,,$ (3)
where $f(x_{t},t)$ and $g(t)$ are respectively the drift and diffusion functions of the PF-ODE defined as follows:
$f(x_{t},t)=\frac{\mathrm{d}\log\alpha(t)}{\mathrm{d}t}x_{t},\hskip 10.00002ptg^{2}(t)=\frac{\mathrm{d}\sigma(t)^{2}}{\mathrm{d}t}-2\frac{\mathrm{d}\log\alpha(t)}{\mathrm{d}t}\sigma^{2}(t)\,.$
$\nabla\log p_{\theta}(x_{t})=-\frac{\varepsilon_{\theta}(x_{t},t)}{\sigma(t)}$ is called the score function of $p_{\theta}(x_{t})$. The PF-ODE can be solved using a neural ODE integrator [7] consisting in iteratively applying the learned function $\varepsilon_{\theta}$ according to given update rules such as the Euler [67] or the Heun solver [23].
A conditional diffusion model can be trained to generate samples from a conditional distribution $p(x_{0}|c)$ by learning conditional denoising functions $\varepsilon_{\theta}(x_{t},t,c)$ or $x_{\theta}(x_{t},t,c)$ [54, 55, 57, 59, 18, 11, 50, 5, 6]. In that particular setting, Classifier-Free Guidance (CFG) [16] has proven to be a very efficient way to better enforce the model to respect the conditioning and so improve the sampling quality. CFG is a technique that consists in dropping the conditioning $c$ with a certain probability during training and replacing the conditional noise estimate $\varepsilon_{\theta}(x_{t},t,c)$ with a linear combination at inference time as follows:
$\varepsilon_{\theta}(x_{t},t,c)=\omega\cdot\varepsilon_{\theta}(x_{t},t,c)+(1-\omega)\cdot\varepsilon_{\theta}(x_{t},t,\varnothing)\,,$ (4)
where $\omega>0$ is called the guidance scale.
3.2 Consistency Models
Since our approach is inspired by the idea exposed in consistency models [68, 42], we recall some elements of those models. Consistency Models (CM) are a new class of generative models designed primarily to learn a consistency function $f_{\theta}$ that maps any sample $x_{t}$ lying on a trajectory of the PF-ODE given in Eq. (3) directly to the original sample $x_{0}$ while ensuring the self-consistency property for any $t\in[\varepsilon,T]$, $\varepsilon>0$ [68, 42, 65]:
$f_{\theta}(x_{t},t)=f_{\theta}(x_{t^{\prime}},t^{\prime}),\hskip 10.00002pt\forall(t,t^{\prime})\in[\varepsilon,T]^{2}\,.$ (5)
In order to ensure the self-consistency property, the authors of [68] proposed to parametrized $f_{\theta}$ as follows:
$f_{\theta}(x_{t},t)=c_{\mathrm{skip}}(t)\cdot x_{t}+c_{\mathrm{out}}(t)\cdot F_{\theta}(x_{t},t)\,,$
where $F_{\theta}$ is parametrized using a neural network and $c_{\mathrm{skip}}$ and $c_{\mathrm{out}}$ are differentiable functions [68, 42]. A consistency model can be trained either from scratch (Consistency Training) or can be used to distill an existing DM (Consistency Distillation) [68, 42]. In both cases, the objective of the model is to learn $f_{\theta}$ such that it matches the output of a target function $f_{\theta^{-}}$ the weights of which are updated using Exponential Moving Average (EMA), for any given points $(x_{t},x_{t^{\prime}})$ lying on a trajectory of the PF-ODE:
$\mathcal{L}=\mathbb{E}_{x_{0}\sim p(x_{0}),t\sim\pi(t),\varepsilon\sim\mathcal{N}\left(0,\mathbf{I}\right)}\left[\left\|f_{\theta}(x_{t},t)-f_{\theta^{-}}(x_{t^{\prime}},t^{\prime})\right\|^{2}\right]\,,$
In other words, given a noisy sample $x_{t}$ obtained with Eq. (1), the idea is to enforce that $f_{\theta}(x_{t},t)=f_{\theta^{-}}(x_{t^{\prime}},t^{\prime})$ where $x_{t^{\prime}}$ is obtained using either Eq. (1) with the same noise $\varepsilon$ and input $x_{0}$ for Consistency Training [68, 65] or using a trained diffusion model $\varepsilon_{\phi}^{\mathrm{teacher}}$ and an ODE solver $\Psi$ for Consistency Distillation [68, 65]. Once the model is trained, one may theoretically generate a sample $\tilde{x}_{0}$ in a single step by first drawing a noisy sample $x_{T}\sim\mathcal{N}\left(0,\mathbf{I}\right)$ and then applying the learned function $f_{\theta}$ to it. In practice, several iterations are required to generate a satisfying sample and so the estimated sample $\tilde{x}_{0}$ is iteratively re-noised and denoised several times using the learned function $f_{\theta}$.
4 Proposed Method
@@
In this section, we expose the proposed method that builds upon several ideas proposed in the literature. For the following, we place ourselves in the context of Latent Diffusion Models [57] for image generation and refer to the teacher model as $\varepsilon_{\phi}^{\mathrm{teacher}}$, the student model as $\varepsilon_{\theta}^{\mathrm{student}}$, the training images as $x_{0}$ and their unknown distribution $p(x_{0})$. We refer to $z_{0}=\mathcal{E}(x_{0})$ as the associated latent variables obtained with an encoder $\mathcal{E}$. We denote $\pi$, the probability density function of the timesteps, and set $T=1$. Note that the presented approach also applies straightforwardly to pixel-space diffusion models.
4.1 Distilling a Pretrained Diffusion Model
@@
The proposed method is mainly driven by the desire to end up with a fast, robust, and reliable approach that would be easily transposed to different use cases. Given a set of data $x_{0}\in\mathcal{X}$ such that $x_{0}\sim p(x_{0})$ and $z_{0}=\mathcal{E}(x_{0})$ the associated latent variables, the main idea of the proposed approach is quite similar to diffusion models. Given a noise schedule defined by $\alpha(t)$ and $\sigma(t)$, we propose to create a noisy latent sample $z_{t}$ with $t\sim\pi(t)$ as specified in Eq. (1) and train a function $f_{\theta}^{\mathrm{student}}$ to predict a denoised version $\tilde{z}_{0}$ of the original sample $z_{0}$. The main difference with a diffusion model is that instead of using $z_{0}$ as a target, we propose to leverage the knowledge of the teacher model and use a sample belonging to the data distribution learned by the teacher model $p_{\phi}^{\mathrm{teacher}}(z_{0})$. In other words, we use the teacher model and an ODE solver to generate a denoised latent sample $\tilde{z}_{0}^{\mathrm{teacher}}(z_{t})$ that belongs to the learned data distribution and use it as a target for the student model. The main distillation loss writes as follows:
@@
$\mathcal{L}_{\mathrm{distil}}=\mathbb{E}_{z_{0},t,\varepsilon}\left[\left\|f_{\theta}^{\mathrm{student}}(z_{t},t)-\tilde{z}_{0}^{\mathrm{teacher}}(z_{t})\right\|^{2}\right]\,,$ (6)
@@
where $\pi(t)$ denotes the distribution over the timesteps, and $\tilde{z}_{0}^{\mathrm{teacher}}(z_{t})$ is obtained by running several steps of an ODE solver $\Psi$ on a teacher model $\varepsilon_{\phi}^{\mathrm{teacher}}$ starting from $z_{t}=\alpha(t)\cdot z_{0}+\sigma(t)\cdot\varepsilon$. A similar idea was employed in [62] but the authors generate fully synthetic samples meaning that the samples $z_{t}$ are pure noise, $z_{t}\sim\mathcal{N}\left(0,\mathbf{I}\right)$. In contrast, in our approach, we hypothesize that allowing $z_{t}$ to retain some information from the ground-truth encoded sample $z_{0}$ could enhance the distillation process. As in [42], when distilling a conditional DM, we also perform Classifier-Free Guidance (CFG) [16] with the teacher to better enforce the model to respect the conditioning. This technique actually significantly improves the quality of the generated samples by the student as shown in Sec. 5. Additionally, it eliminates the need for conducting CFG during inference with the student, further decreasing the method’s computational cost by halving the NFEs for each step. The value of the guidance scale $\omega$ used during training is part of the ablations presented in Sec. 5.2 but in practice $\omega$ is uniformly sampled in $[\omega_{\min},\omega_{\max}]$ where $0\leq\omega_{\min}\leq\omega_{\max}$. As mentioned in Sec. 3.2, our approach bears resemblance to existing consistency models [68, 42]. However, rather than relying on a previous instance of the student model to estimate the origin of the PF-ODE, we directly employ the teacher model coupled with an ODE solver to generate the target. We observed that these ingredients enhance the stability of the training procedure.
@@
4.2 Timesteps Sampling
The cornerstone of our approach hinges on the selection of the timestep probability density function, denoted as $\pi(t)$. According to the continuous modeling, exposed in [67], DMs are trained to remove noise from a latent sample $z_{t}$ for any given continuous time $t$. However, since we aim at achieving few steps data generation (typically 1-4 steps) at inference time, the learned function $\varepsilon_{\theta}$ will only be evaluated at a few discrete timesteps $\{t_{i}\}_{i=1}^{K}$.
In order to tackle this issue and enforce the distillation process to focus on the most relevant timesteps, we propose to select $K$ (typically 16, 32, or 64) uniformly spaced timesteps in the range $[0,1]$ and assign a probability to each of them according to a probability mass function $\pi(t)$. We choose $\pi(t)$ as a mixture of Gaussian distribution controlled by a series of weights $\{\beta_{i}\}_{i=1}^{K}$

### v3

@@
Flash Diffusion: Accelerating Any Conditional Diffusion Model for Few Steps Image Generation
Clément Chadebec, Onur Tasar , Eyal Benaroche, Benjamin Aubin Corresponding author: [name].[surname]@jasper.aiWork done during an internship at Jasper Research
In this paper, we propose an efficient, fast, and versatile distillation method to accelerate the generation of pre-trained diffusion models. The method reaches state-of-the-art performances in terms of FID and CLIP-Score for few steps image generation on the COCO2014 and COCO2017 datasets, while requiring only several GPU hours of training and fewer trainable parameters than existing methods. In addition to its efficiency, the versatility of the method is also exposed across several tasks such as text-to-image, inpainting, face-swapping, super-resolution and using different backbones such as UNet-based denoisers (SD1.5, SDXL), DiT (Pixart-$\alpha$) and MMDiT (SD3), as well as adapters. In all cases, the method allowed to reduce drastically the number of sampling steps while maintaining very high-quality image generation.
Codehttps://github.com/gojasper/flash-diffusion
@@ Introduction
Diffusion Models (DM) (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020; Song et al. 2020) have proven to be one of the most efficient class of generative models for image synthesis (Dhariwal and Nichol 2021; Ramesh et al. 2022; Rombach et al. 2022; Nichol et al. 2022) and have raised particular interest and enthusiasm for text-to-image applications (Ramesh et al. 2021, 2022; Rombach et al. 2022; Saharia et al. 2022; Ho et al. 2022; Esser et al. 2024; Podell et al. 2023; Chen et al. 2023, 2024) where they outperform other approaches. However, their usability for real-time applications remains limited by the intrinsic iterative nature of their sampling mechanism. At inference time, these models aim at iteratively denoising a sample drawn from a Gaussian distribution to finally create a sample belonging to the data distribution. Nonetheless, such a denoising process requires multiple evaluations of a potentially very computationally costly neural function.
@@
Recently, more efficient solvers (Lu et al. 2022a, b; Zhang and Chen 2022; Zhao et al. 2024) or diffusion distillation methods (Salimans and Ho 2021; Song et al. 2023; Lin, Wang, and Yang 2024; Xu et al. 2023; Liu et al. 2023; Ren et al. 2024; Luo et al. 2023a, b; Sauer et al. 2023, 2024; Yin et al. 2023; Hsiao et al. 2024) aiming at reducing the number of sampling steps required to generate satisfying samples from a trained diffusion model have emerged to try to tackle this issue. Nonetheless, solvers typically require at least 10 Neural Function Evaluations (NFEs) to produce satisfying samples while distillation methods may require extensive training resources (Liu et al. 2023; Yin et al. 2023; Meng et al. 2023) or require an iterative training procedure to update the teacher model throughout training (Salimans and Ho 2021; Lin, Wang, and Yang 2024; Li et al. 2024) limiting their applications and reach. Moreover, most of the existing distillation methods are tailored for a specific task such as text-to-image. It is still unclear how they would perform on other tasks, using different conditionings and diffusion model architectures.
@@
In this paper, we present Flash Diffusion, a fast, robust, and versatile diffusion distillation method that allows to drastically reduce the number of sampling steps while maintaining a very high image generation quality. The proposed method aims at training a student model to predict in a single step a denoised multiple-step teacher prediction of a corrupted input sample. The method also drives the student distribution towards the real input sample manifold with an adversarial objective (Goodfellow et al. 2014) and ensures that it does not drift too much from the learned teacher distribution using distribution matching (Dziugaite, Roy, and Ghahramani 2015; Li, Swersky, and Zemel 2015). The main contributions of the paper are as follows:
@@
  • We validate the method for text-to-image and show that it reaches SOTA results for few steps image generation on standard benchmark datasets with only two NFEs, which is equivalent to a single step with classifier-free guidance while having far fewer training parameters than competitors and requiring only a few GPU hours of training.
@@
  • We emphasize the versatility of the method through an extensive experimental study across various tasks (text-to-image, image inpainting, super-resolution, face-swapping), diffusion model architectures (SD1.5, SDXL, Pixart-$\alpha$ and SD3) and illustrate its compatibility with adapters (Mou et al. 2024) and existing LoRAs.
Related Works
@@
Diffusion models consist in artificially corrupting input data according to a given noise schedule (Sohl-Dickstein et al. 2015; Ho, Jain, and Abbeel 2020; Song et al. 2020) such that the data distribution eventually resembles a standard Gaussian one. They are then trained to estimate the amount of noise added in order to learn a reverse diffusion process allowing them, once trained, to generate new samples from Gaussian noise. Those models can be conditioned with respect to various inputs such as images (Rombach et al. 2022), depth maps, edges, poses (Zhang, Rao, and Agrawala 2023; Mou et al. 2024) or text (Dhariwal and Nichol 2021; Ramesh et al. 2022; Rombach et al. 2022; Nichol et al. 2022; Esser et al. 2024; Ho et al. 2022; Podell et al. 2023) where they demonstrated very impressive results. However, the need to recourse to a large number of sampling steps (typically 50 steps) at inference time to generate high-quality samples has limited their usage for real-time applications and narrowed their usability and reach.
@@
In order to tackle this limitation, several methods have recently emerged to reduce the number of function evaluations required at inference time. On the one hand, several papers tried to build more efficient solvers to speed up the generation process (Lu et al. 2022a, b; Zhang and Chen 2022; Zhao et al. 2024) but these methods still require the use of several steps (typically 10) to generate satisfying samples. On the other hand, several approaches relying on model distillation (Hinton, Vinyals, and Dean 2015) proposed to train a student network that would learn to match the samples generated by a teacher model but in fewer steps. A simple approach would consist in building pairs of noise/teacher samples and training a student model to match the teacher predictions in a single step (Luhman and Luhman 2021; Zheng et al. 2023). Nonetheless, this approach remains quite limited and struggles to match the quality of the teacher model since there is no underlying useful information to be learned by the student in full noise. Building upon this idea, several methods were proposed to first apply the forward diffusion process to an input sample and then pass it to the student network. The student prediction is then compared to the learned distribution of the teacher model using either a regression loss (Kohler et al. 2024; Yin et al. 2023) an adversarial objective (Xu et al. 2023; Sauer et al. 2023, 2024; Yin et al. 2024) or distribution matching (Yin et al. 2023, 2024).
@@
Progressive distillation (Salimans and Ho 2021; Meng et al. 2023) is also a method that has proven to be quite promising. It consists in training a student model to predict a two-step teacher denoising of a noisy sample in a single step theoretically halving the number of required sampling steps. The teacher is then replaced by the new student and the process is repeated several times. This approach was also enriched with a GAN-based objective that allows to further reduce the number of sampling steps needed from 4-8 to a single pass (Lin, Wang, and Yang 2024). InstaFlow (Liu et al. 2023) proposed instead to rely on rectified flows (Liu, Gong et al. 2022) to ease the one-step distillation process. However, this approach may require a significant number of training parameters and a long training procedure, making it computationally intensive.
@@
Consistency models (Song et al. 2023; Song and Dhariwal 2023; Luo et al. 2023a; Kim et al. 2023) is also a promising, effective, and one of the most versatile distillation methods proposed in the literature. The main idea is to train a model to map any point lying on the Probability Flow Ordinary Differential Equation (PF-ODE) to its origin, theoretically unlocking single-step generation. Luo et al. (2023b) combined Latent Consistency Model (LCM) and LoRAs (Hu et al. 2021) and showed that it is possible to train a strong student with a very limited number of trainable parameters and a few GPU hours of training. Nonetheless, those models still struggle to achieve single-step generation and reach the sampling quality of peers.
@@
In a parallel study conducted recently, the authors of (Yin et al. 2024) also introduced the combined use of a distribution matching loss and an adversarial loss, a method we also employ in our paper. Nonetheless, they do not rely on the use of a distillation loss that proved highly efficient in our experiments and do not compute the adversarial loss with respect to the same inputs. Moreover, their approach still necessitates training another denoiser to assess the score of the fake samples, significantly increasing the number of trainable parameters and, consequently, the computational burden of the method. Furthermore, the ability of their method to generalize and perform effectively across different tasks and diffusion model architectures remains unclear.
Proposed Method
In this section, we expose the proposed method that builds upon several ideas proposed in the literature.
Background on Diffusion Models
Let $x_{0}\in\mathcal{X}$ be a set of data such that $x_{0}\sim p(x_{0})$ where $p(x_{0})$ is an unknown distribution. The main idea of diffusion models (DM) is to estimate the amount of noise $\varepsilon$, artificially added to an input sample $x_{0}$ using the forward process $x_{t}=\alpha(t)\cdot x_{0}+\sigma(t)\cdot\varepsilon$ where $\varepsilon\sim\mathcal{N}\left(0,\mathbf{I}\right)$. The noise schedule is controlled by two differentiable functions $\alpha(t)$, $\sigma(t)$ for any $t\in[0,T]$ such that the log signal-to-noise ratio $\log[\alpha(t)^{2}/\sigma(t)^{2}]$ is decreasing over time. In practice, during training a diffusion model learns a parametrized function $\varepsilon_{\theta}$ conditioned on the timestep $t$ and taking as input the noisy sample $x_{t}$. The parameters $\theta$ are then learned via denoising score matching (Vincent 2011; Song and Ermon 2019).
@@
$\mathcal{L}=\mathbb{E}_{x_{0}\sim p,t\sim\pi,\varepsilon\sim\mathcal{N}\left(0,\mathbf{I}\right)}\left[\lambda(t)\left\|\varepsilon_{\theta}(x_{t},t)-\varepsilon\right\|^{2}\right]\,,$ (1)
where $\lambda(t)$ is a scaling factor, $t\in[0,1]$ is the timestep and $\pi(t)$ is a distribution over the timesteps. We provide in the appendices an extended background on diffusion models.
Distilling a Pretrained Diffusion Model
@@
For the following, we place ourselves in the context of Latent Diffusion Models (Rombach et al. 2022) for image generation and refer to the teacher model as $\varepsilon_{\phi}^{\mathrm{teacher}}$, the student model as $\varepsilon_{\theta}^{\mathrm{student}}$, the training images as $x_{0}$ and their unknown distribution $p(x_{0})$. We refer to $z_{0}=\mathcal{E}(x_{0})$ as the associated latent variables obtained with an encoder $\mathcal{E}$. $\pi$ is the probability density function of the timesteps $t\in[0,1]$. The proposed method is mainly driven by the desire to end up with a fast, robust, and reliable approach that would be easily transposed to different use cases. The main idea of the proposed approach is quite similar to diffusion models.
@@
Given a noisy latent sample $z_{t}$ with $t\sim\pi(t)$, we propose to train a function $f_{\theta}$ to predict a denoised version $\tilde{z}_{0}$ of the original sample $z_{0}$. The main difference with a diffusion model is that instead of using $z_{0}$ as a target, we propose to leverage the knowledge of the teacher model and use a sample belonging to the data distribution learned by the teacher model $p_{\phi}^{\mathrm{teacher}}(z_{0})$. In other words, we use the teacher model and an ODE solver $\Psi$ that is run several times to generate a denoised latent sample $\tilde{z}_{0}^{\mathrm{teacher}}(z_{t})$ used as a target for the student model. The main distillation loss writes as follows:
@@
$\mathcal{L}_{\mathrm{distil}}=\mathbb{E}_{z_{0},t,\varepsilon}\left[\left\|f_{\theta}(z_{t},t)-\tilde{z}_{0}^{\mathrm{teacher}}(z_{t})\right\|^{2}\right]\,,$ (2)
@@
A similar idea was employed in (Sauer et al. 2024) but the authors generate fully synthetic samples meaning that the samples $z_{t}$ are pure noise, $z_{t}\sim\mathcal{N}\left(0,\mathbf{I}\right)$. In contrast, in our approach, we hypothesize that allowing $z_{t}$ to retain some information from the ground-truth encoded sample $z_{0}$ could enhance the distillation process. As in (Luo et al. 2023a), when distilling a conditional DM, we also perform Classifier-Free Guidance (CFG) (Ho and Salimans 2021) with the teacher to better enforce the model to respect the conditioning. This technique actually significantly improves the quality of the generated samples by the student as shown in the ablations. Additionally, it eliminates the need for conducting CFG during inference with the student, further decreasing the method’s computational cost by halving the NFEs for each step. In practice, the guidance scale $\omega$ is uniformly sampled in $[\omega_{\min},\omega_{\max}]$ where $0\leq\omega_{\min}\leq\omega_{\max}$.
@@
Timesteps Sampling
The cornerstone of our approach hinges on the selection of the timestep probability density function, denoted as $\pi(t)$. According to the continuous modeling, exposed in (Song et al. 2020), DMs are trained to remove noise from a latent sample $z_{t}$ for any given continuous time $t$. However, since we aim at achieving few steps data generation (typically 1-4 steps) at inference time, the learned function $\varepsilon_{\theta}$ will only be evaluated at a few discrete timesteps $\{t_{i}\}_{i=1}^{K}$.
To tackle this issue and enforce the distillation process to focus on the most relevant timesteps, we propose to select $K$ (typically 16 or 32) uniformly spaced timesteps in $[0,1]$ and assign a probability to each of them according to a probability mass function $\pi(t)$. We choose $\pi(t)$ as a mixture of Gaussian controlled by a series of weights $\{\beta_{i}\}_{i=1}^{K}$